Joker stock has a sustainable growth rate of 9 percent, ROE of 17 percent, and dividends per share of $1.55. If the P/E ratio is 15.5, what is the value of a share of stock? (Do not round intermediate calculations. Round your answer to 2 decimal places. Omit the "$" sign in your response.)

Share of stock $______
Answer #1

Sustainable growth rate = 9%

ROE = 17%

Dividend = $1.55

Sustainable growth rate = ROE x retention rate

Sustainable growth rate = ROE x (1- Dividend payout ratio)

Sustainable growth rate = ROE x (1- DPS / EPS)

9% = 17% x (1-DPS / EPS)

0.529412 = 1 - DPS/EPS

DPS / EPS = 1 - 0.529412

1.55 / EPS = 0.470588

EPS = 3.29375

P/E = 15.5

Price / Earning per share = 15.5

Price = 15.5 x 3.29375 = 49.41